ADNAN MUHAMMED

Grade 12,

Calculate the mass of a photon with wavelength 3.6 Å

Calculate the mass of a photon with wavelength 3.6 Å

Grade:12

10 Answers

Muneera
8 years ago
Lambda=3.6Å=3.6*10-10m 
Velocity of photon =velocity of light
M=h/lambda*v =6.626*10-34js/(3.6*10-10m)(3*10 8 m/s)
=6.135*10 -29 Kg
Prachi
7 years ago
lamda = 3.6 angstrom =  3.6 x 10^-10 m
velocity of photon = velocity of light 
m = h/lamda*v ( from de-brogile’s equation)
m = 6.626 x 10^-34/3.6 x10^-10 x 3 x 10^8
m = 6.626 x 10^-32/10.8
m = 0.6135 x 10^-32
m = 6.135 x 10^-33 kg
 
Udayveer
6 years ago
We know that wavelength=Planck`s constant/mass ×velocityOr lambda=h/mvm=h/lambda×v=6.63×10^-34 / 3.6×10^-10 × 3×10^8=6.139 × 10^-33
Rahul chouhan
6 years ago
We know that wavelength=Planck`s constant/mass ×velocityOr lambda=h/mvm=h/lambda×v=6.63×10^-34 / 3.6×10^-10 × 3×10^8=6.139 × 10^-33
Perunchelvan
6 years ago
Mass of photon m=h÷lambda×c. This is derived from debrogle equation. m=6.626×10^-34÷3.6×3×10^8×10^-10=0.6135×10^-32 kg m
Harshit Panwar
6 years ago
Mas
shibani
6 years ago
we know that  wavelength=planks constant/mass*velocity. from that we calculate mass=6.26*10^-34/3.6*3*10^8*10^-10
Ashish D Parbale
6 years ago
H/P=frequency6.626×10^-34/p=3.6×10^-10P=6.626×10^_34/3.6×10^-10P=1.8×10^-24P=MV1.8×10^-24=m×3×10^8M=6×10^-33
Venkata poojitha
5 years ago
We know that wave length=plancks constant/mass *velocity
 
Here velocity is speed of light=3*10
3.6*10
 
By calculating this we can get the mass 
 
Answer is 6.135*10
 
 
 
Krish Gupta
3 years ago

Momentum=p

p=h/lambda=mass*velocity

mass=h/(lambda*velocity)

mass=6.63*10^(-34)/(3. 6*10^(-10)*3*10^(8))

mass=0.613888*10^(-32)kg

mass=6.13888*10^(-33)kg

 

Cheers !